10. Benzene ( C6H6 ) and toluene ( C6H5CH3 ) form ideal solutions. At 35°C the vapor pressure of benzene is 160. torr and that of toluene is 50.0 torr. If 3.6 mol of benzene and 5.5 mol of toluene are placed in a closed container at 35°C, what is the mole fraction of toluene in the vapor phase when equilibrium is reached? A) 0.52 B) 0.68 C) 0.32 D) 0.14 E) 0.60
Answer #1

X(benzene, ℓ)(160 torr) + X(toluene, ℓ)(50 torr)
= (3.6/9.1)(160 torr) + (5.5/9.1)(50 torr) = 63.297 + 30.22 = 93.52 torr = P(tot)

X(toluene, v) = 30.22/93.52 = 0.323 ⇐ answer is C).